OpenStudy (anonymous):

how to find domain of real-valued function ( g o h) g(x) 4x-1 h(x) sq rt (x-4) close sq rt

OpenStudy (anonymous):

interval notation

OpenStudy (anonymous):

is the question h(x)=sqrt(x-4) ? or some thing more with h(x)?

OpenStudy (anonymous):

\[(g\circ h)(x) = g(h(x)) = 4(\sqrt{x-4}) - 1\]

OpenStudy (anonymous):

You just plug in h(x) for x in g(x)

OpenStudy (anonymous):

i got that much but what is the domain ??

OpenStudy (anonymous):

so here for the square root to be defined x-4>=0 and hence x>=4 therefore the domain is x>=4

OpenStudy (anonymous):

so in interval notation (4, \[\infty\]

OpenStudy (anonymous):

whoops lol (4, infinity]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[x \in [4,\infty)\]

OpenStudy (anonymous):

x can be 4, and any other number up to (but never including) infinity.

OpenStudy (anonymous):

thnks